# Deep learning reveals key predictors of thermal conductivity in covalent organic frameworks

### v2

@@ Discussion
Further analysis using the transformer’s attention mechanism revealed that COFs with higher amounts of dangling atoms exhibited lower thermal conductivities due to disrupted heat transfer pathways, identifying a novel and significant predictor of thermal conductivity. The random forest regressor, identified as the best-performing regression ensemble model, was analyzed using Gini importance, permutation feature importance, and SHAP values. These analyses consistently highlighted the dangling mass ratio as the second most important predictor of thermal conductivity. The vibrational density of states analysis further supported these findings, showing that dangling masses introduce mismatched vibrational modes that hinder thermal transfer.
@@ Methods
The molecular dynamics setup, the distribution of in-plane thermal conductivities, and feature correlations (including DMR) (Section 1); implementation details of the ensemble regression models and regression performances excluding DMR (Section 2); the deep learning setup (Section 3); the dangling mass computation (Section 4); additional comparisons of COFs with and without dangling mass, including phonon dispersion maps for all examples (Section 5); details on the VDOS and pSED computation (Section 6); and estimated runtimes (Section 7) are summarized in the Supporting Information.
@@ Acknowledgements
Competing interests
The authors declare no competing interests.
Data availability
The datasets generated during the current study are available at <link will be provided at time of publication>.
Code availability
The codes generated during the current study are available at <link will be provided at time of publication>.

### v3

@@ Discussion
Further analysis using the transformer’s attention mechanism revealed that COFs with higher amounts of dangling atoms exhibited lower thermal conductivities due to disrupted heat transfer pathways, identifying a novel and significant predictor of thermal conductivity. The random forest regressor, identified as the best-performing regression ensemble model, was used primarily as a simple tool to assess feature importance rather than to compare predictive performance with the PMTransformer. It was analyzed using Gini importance, permutation feature importance, and SHAP values. These analyses consistently highlighted the dangling mass ratio as the second most important predictor of thermal conductivity. The vibrational density of states analysis further supported these findings, showing that dangling masses introduce mismatched vibrational modes that hinder thermal transfer. While our current analysis focuses on two-dimensional monolayer COFs, additional mechanisms may emerge in three-dimensional architectures and warrant separate investigation.
@@ Methods
The molecular dynamics setup, the distribution of in-plane thermal conductivities, and feature correlations (including DMR) (Section 1); implementation details of the ensemble regression models and regression performances excluding DMR (Section 2); the deep learning setup (Section 3); the dangling mass computation (Section 4); additional comparisons of COFs with and without dangling mass, including phonon dispersion maps for all examples (Section 5); details on the VDOS and pSED computation (Section 6); and estimated runtimes (Section 7) are summarized in the Supplementary Information.
Author contributions
P.T and Y.Z.: conceptualization, data curation, formal analysis, investigation, writing – original draft. P.T., Y.Z., Y.G.: methodology, software, validation, visualization. A.V. and S.K.: conceptualization, supervision, funding acquisition, writing – review & editing
Competing interests
The authors declare no competing interests.
Data and code availability
The code and data generated during this study are available in the following repository: doi.org/10.4121/5866cc9a-78bf-4a0c-9280-ae526da86ac9. The code and data are also available at: github.com/mmc-group/deep-learning-thermal-conductivity-of-COFs. The data includes unit-cell CIF files obtained from the dataset of Mercado et al.47 available at doi.org/10.24435/materialscloud:2018.0003/v2. We also used the dataset provided by Lan et al.65, which is available at figshare.com/s/c7e3b7610a71b9d64210 with DOI: 10.1038/s41467-018-07720-x for the associated paper. The codes are built upon the open-source MOFTransformer/PMTransformer repository by Park et al.46, which can be accessed at github.com/hspark1212/MOFTransformer with DOI: 10.1021/acsami.3c10323 for the associated paper.
@@ Acknowledgements
Notes and references
  • Wang et al. 2020 H. Wang, H. Wang, Z. Wang, L. Tang, G. Zeng, P. Xu, M. Chen, T. Xiong, C. Zhou, X. Li, D. Huang, Y. Zhu, Z. Wang and J. Tang, Chemical Society Reviews, 2020, 49, 4135–4165.
  • Zhao et al. 2021 X. Zhao, P. Pachfule and A. Thomas, Chemical Society Reviews, 2021, 50, 6871–6913.
  • Liu et al. 2021 C. Liu, J. Wang, J. Wan and C. Yu, Coordination Chemistry Reviews, 2021, 432, 213743.
  • Stock and Biswas 2011 N. Stock and S. Biswas, Chemical Reviews, 2011, 112, 933–969.
  • Pérez-Botella et al. 2022 E. Pérez-Botella, S. Valencia and F. Rey, Chemical Reviews, 2022, 122, 17647–17695.
  • Abuzeid et al. 2021 H. R. Abuzeid, A. F. EL-Mahdy and S.-W. Kuo, Giant, 2021, 6, 100054.
  • Dogru et al. 2013 M. Dogru, M. Handloser, F. Auras, T. Kunz, D. Medina, A. Hartschuh, P. Knochel and T. Bein, Angewandte Chemie International Edition, 2013, 52, 2920–2924.
  • Stegbauer et al. 2018 L. Stegbauer, S. Zech, G. Savasci, T. Banerjee, F. Podjaski, K. Schwinghammer, C. Ochsenfeld and B. V. Lotsch, Advanced Energy Materials, 2018, 8, 1703278.
  • Ali et al. 2024 Z. Ali, T. Huo, Y. Zhang and G. Wang, Microchemical Journal, 2024, 200, 110340.
  • Li et al. 2020 Y. Li, M. Chen, Y. Han, Y. Feng, Z. Zhang and B. Zhang, Chemistry of Materials, 2020, 32, 2532–2540.
  • Yusran et al. 2020 Y. Yusran, H. Li, X. Guan, Q. Fang and S. Qiu, EnergyChem, 2020, 2, 100035.
  • Scicluna and Vella-Zarb 2020 M. C. Scicluna and L. Vella-Zarb, ACS Applied Nano Materials, 2020, 3, 3097–3115.
  • Guo et al. 2022 H. Guo, Y. Liu, N. Wu, L. Sun and W. Yang, ChemistrySelect, 2022, 7, e202202538.
  • Wang et al. 2017 L. Wang, B. Dong, R. Ge, F. Jiang and J. Xu, ACS Applied Materials & Interfaces, 2017, 9, 7108–7114.
  • Chumakov et al. 2016 Y. Chumakov, F. Aksakal, A. Dimoglo, A. Ata and S. A. Palomares-Sánchez, Journal of Electronic Materials, 2016, 45, 3445–3452.
  • Wang et al. 2020 S. Wang, X. Xu, Y. Yue, K. Yu, Q. Shui, N. Huang and H. Chen, Small Structures, 2020, 1, 2000021.
  • Fan et al. 2018 H. Fan, A. Mundstock, A. Feldhoff, A. Knebel, J. Gu, H. Meng and J. Caro, Journal of the American Chemical Society, 2018, 140, 10094–10098.
  • Wu and Yang 2017 M.-X. Wu and Y.-W. Yang, Chinese Chemical Letters, 2017, 28, 1135–1143.
  • Zhou et al. 2018 X. Zhou, Y. Yan, X. Lu, H. Zhu, X. Han, G. Chen and Z. Ren, Materials Today, 2018, 21, 974–988.
  • Evans et al. 2021 A. M. Evans, M. R. Ryder, W. Ji, M. J. Strauss, A. R. Corcos, E. Vitaku, N. C. Flanders, R. P. Bisbey and W. R. Dichtel, Faraday Discussions, 2021, 225, 226–240.
  • Zou et al. 2013 X. Zou, H. Ren and G. Zhu, Chemical Communications, 2013, 49, 3925.
  • Pang et al. 2017 Z.-F. Pang, T.-Y. Zhou, R.-R. Liang, Q.-Y. Qi and X. Zhao, Chemical Science, 2017, 8, 3866–3870.
  • Gui et al. 2020 B. Gui, G. Lin, H. Ding, C. Gao, A. Mal and C. Wang, Accounts of Chemical Research, 2020, 53, 2225–2234.
  • Xu et al. 2018 X. Xu, J. Chen, J. Zhou and B. Li, Advanced Materials, 2018, 30, 1705544.
  • Choy 1977 C. Choy, Polymer, 1977, 18, 984–1004.
  • Utimula et al. 2019 K. Utimula, T. Ichibha, R. Maezono and K. Hongo, Chemistry of Materials, 2019, 31, 4649–4656.
  • Lindsay et al. 2012 L. Lindsay, D. A. Broido and T. L. Reinecke, Phys. Rev. Lett., 2012, 109, 095901.
  • Romero et al. 2015 A. H. Romero, E. K. U. Gross, M. J. Verstraete and O. Hellman, Phys. Rev. B, 2015, 91, 214310.
  • Li and Brédas 2021 H. Li and J.-L. Brédas, Chemistry of Materials, 2021, 33, 4529–4540.
  • Li et al. 2020 Z. Li, Z. Liu, Z. Li, T. Wang, F. Zhao, X. Ding, W. Feng and B. Han, Advanced Functional Materials, 2020, 30, 1909267.
  • Rapaport 2004 D. C. Rapaport, The Art of Molecular Dynamics Simulation, Cambridge University Press, 2004.
  • Islamov et al. 2023 M. Islamov, H. Babaei, R. Anderson, K. B. Sezginel, J. R. Long, A. J. H. McGaughey, D. A. Gomez-Gualdron and C. E. Wilmer, npj Computational Materials, 2023, 9, DOI: 10.1038/s41524–022–00961–x.
  • Thakur and Giri 2024 S. Thakur and A. Giri, Small, 2024, 20, 2401702.
  • Huang et al. 2023 X. Huang, S. Ma, C. Y. Zhao, H. Wang and S. Ju, npj Computational Materials, 2023, 9, DOI: 10.1038/s41524–023–01154–w.
  • Ma et al. 2022 R. Ma, H. Zhang and T. Luo, ACS Applied Materials & Interfaces, 2022, 14, 15587–15598.
  • Wu et al. 2019 S. Wu, Y. Kondo, M.-a. Kakimoto, B. Yang, H. Yamada, I. Kuwajima, G. Lambard, K. Hongo, Y. Xu, J. Shiomi et al., Npj Computational Materials, 2019, 5, 66.
  • Ishikiriyama 2022 K. Ishikiriyama, Thermochimica Acta, 2022, 708, 179135.
  • Xu and Luo 2024 J. Xu and T. Luo, npj Computational Materials, 2024, 10, 74.
  • Hu et al. 2024 Y. Hu, Q. Wang and H. Ma, Journal of Applied Physics, 2024, 135, 120701.
  • Yang et al. 2021 P. Yang, H. Zhang, X. Lai, K. Wang, Q. Yang and D. Yu, ACS Omega, 2021, 6, 17149–17161.
  • Cao et al. 2022 X. Cao, Z. Zhang, Y. He, W. Xue, H. Huang and C. Zhong, Industrial & Engineering Chemistry Research, 2022, 61, 11116–11123.
  • De Vos et al. 2024 J. S. De Vos, S. Ravichandran, S. Borgmans, L. Vanduyfhuys, P. Van Der Voort, S. M. J. Rogge and V. Van Speybroeck, Chemistry of Materials, 2024, 36, 4315–4330.
  • Zhao and Chung 2024 G. Zhao and Y. G. Chung, Journal of Chemical Theory and Computation, 2024, 20, 5368–5380.
  • Korolev and Mitrofanov 2024 V. Korolev and A. Mitrofanov, Journal of Chemical Information and Modeling, 2024, 64, 1919–1931.
  • Kang et al. 2023 Y. Kang, H. Park, B. Smit and J. Kim, Nature Machine Intelligence, 2023, 5, 309–318.
  • Park et al. 2023 H. Park, Y. Kang and J. Kim, ACS Applied Materials & Interfaces, 2023, 15, 56375–56385.
  • Mercado et al. 2018 R. Mercado, R.-S. Fu, A. V. Yakutovich, L. Talirz, M. Haranczyk and B. Smit, Chemistry of Materials, 2018, 30, 5069–5086.
  • Freitas et al. 2017 S. K. S. Freitas, R. S. Borges, C. Merlini, G. M. O. Barra and P. M. Esteves, The Journal of Physical Chemistry C, 2017, 121, 27247–27252.
  • Vaswani et al. 2017 A. Vaswani, N. Shazeer, N. Parmar, J. Uszkoreit, L. Jones, A. N. Gomez, L. Kaiser and I. Polosukhin, arXiv, 2017, preprint, arXiv:1706.03762, https://arxiv.org/abs/1706.03762.
  • Xie and Grossman 2018 T. Xie and J. C. Grossman, Phys. Rev. Lett., 2018, 120, 145301.
  • 51 GitHub - Sangwon91/GRIDAY: Energy shape calculator for the porous materials — github.com, https://github.com/Sangwon91/GRIDAY, [Accessed 14-06-2024].
  • Dosovitskiy et al. 2020 A. Dosovitskiy, L. Beyer, A. Kolesnikov, D. Weissenborn, X. Zhai, T. Unterthiner, M. Dehghani, M. Minderer, G. Heigold, S. Gelly, J. Uszkoreit and N. Houlsby, arXiv, 2020, prepint, arXiv:2010.11929v2, https://arxiv.org/abs/2010.11929.
  • Luo et al. 2017 D. Luo, C. Huang and Z. Huang, Journal of Heat Transfer, 2017, 140, 031302.
  • Feng et al. 2020 T. Feng, J. He, A. Rai, D. Hun, J. Liu and S. S. Shrestha, Phys. Rev. Appl., 2020, 14, 044023.
  • Feng et al. 2020 D. Feng, Y. Feng, Y. Liu, W. Zhang, Y. Yan and X. Zhang, The Journal of Physical Chemistry C, 2020, 124, 8386–8393.
  • Ying et al. 2021 P. Ying, J. Zhang and Z. Zhong, The Journal of Physical Chemistry C, 2021, 125, 12991–13001.
  • Wang et al. 2023 S. Wang, L. Ren, M. Han, W. Zhou, C. Wong, X. Bai, R. Sun and X. Zeng, Nanoscale, 2023, 15, 8706–8715.
  • Ying et al. 2020 P. Ying, J. Zhang, X. Zhang and Z. Zhong, The Journal of Physical Chemistry C, 2020, 124, 6274–6283.
  • Breiman 2001 L. Breiman, Machine Learning, 2001, 45, 5–32.
  • Friedman 2001 J. H. Friedman, The Annals of Statistics, 2001, 29, 1189–1232.
  • Chen and Guestrin 2016 T. Chen and C. Guestrin, Proceedings of the 22nd ACM SIGKDD International Conference on Knowledge Discovery and Data Mining, 2016, p. 785–794.
  • Freund and Schapire 1997 Y. Freund and R. E. Schapire, Journal of Computer and System Sciences, 1997, 55, 119–139.
  • Pedregosa et al. 2011 F. Pedregosa, G. Varoquaux, A. Gramfort, V. Michel, B. Thirion, O. Grisel, M. Blondel, P. Prettenhofer, R. Weiss, V. Dubourg, J. Vanderplas, A. Passos, D. Cournapeau, M. Brucher, M. Perrot and E. Duchesnay, Journal of Machine Learning Research, 2011, 12, 2825–2830.
  • Lundberg 2017 S. Lundberg, arXiv, 2017, preprint, arXiv:1705.07874v2, https://arxiv.org/abs/1705.07874.
  • Lan et al. 2018 Y. Lan, X. Han, M. Tong, H. Huang, Q. Yang, D. Liu, X. Zhao and C. Zhong, Nature Communications, 2018, 9, 5274.